A box contains 18 coins of nickels and quarters. How many quarters and nickels must be in the box so that it contains $3.30? I have to write an equation first. Thanks
Answer by mananth(16946)


A box contains 18 coins of nickels and quarters.
let n be the number of nickels
q be the number of quarters
n+q=18..............................(1)
5n+25q=330
/5
n+5q=66...............................(2)
subtract the equations
4q=48
q=12
So
n= 6